explain clearly.



∫(7𝑐𝑠𝑐2𝑥 + 2 sec 𝑥 tan 𝑥)𝑑𝑥

1
Expert's answer
2022-05-31T09:58:45-0400
"\int(7\csc2x+2 \sec x \tan x)dx"

"=\int(\dfrac{7}{2\sin x\cos x}+\dfrac{2\sin x}{\cos^2x})dx"

"=\int(\dfrac{7\sin x}{2\sin^2 x\cos x}+\dfrac{2\sin x}{\cos^2x})dx"

Use "u" -substitution


"u=\cos x, du=-\sin x dx"

"\int(\dfrac{7\sin x}{2\sin^2 x\cos x}+\dfrac{2\sin x}{\cos^2x})dx"

"=\int(-\dfrac{7}{2u(1-u^2)}-\dfrac{2}{u^2})du"

"-\dfrac{7}{2u(1-u^2)}=\dfrac{A}{u}+\dfrac{B}{1-u}+\dfrac{C}{1+u}"

"=\dfrac{A(1-u^2)+Bu(1+u)+Cu(1-u)}{u(1-u^2)}"


"-\dfrac{7}{2}=A(1-u^2)+Bu(1+u)+Cu(1-u)"


"u=0:A=-\dfrac{7}{2}"

"u=-1:C=\dfrac{7}{4}"

"u=1:B=-\dfrac{7}{4}"

"\int(-\dfrac{7}{2u(1-u^2)})du"

"=-\dfrac{7}{2}\int\dfrac{du}{u}-\dfrac{7}{4}\int\dfrac{du}{1-u}+\dfrac{7}{4}\int\dfrac{du}{1+u}"

"=-\dfrac{7}{2}\ln |u|+\dfrac{7}{4}\ln|1-u|+\dfrac{7}{4}\ln|1+u|+C_1"

"\int(-\dfrac{2}{u^2})du=\dfrac{2}{u}+C_2"

Then


"\int(-\dfrac{7}{2u(1-u^2)}-\dfrac{2}{u^2})du"

"=-\dfrac{7}{2}\ln |u|+\dfrac{7}{4}\ln|1-u|+\dfrac{7}{4}\ln|1+u|+\dfrac{2}{u}+C"

So


"\int(7\csc2x+2 \sec x \tan x)dx"

"=-\dfrac{7}{2}\ln |\cos x|+\dfrac{7}{4}\ln|1-\cos x|+\dfrac{7}{4}\ln|1+\cos x|"

"+\dfrac{2}{\cos x}+C"

"=-\dfrac{7}{2}\ln |\cos x|+\dfrac{7}{4}\ln|1-\cos^2 x|+\dfrac{2}{\cos x}+C"

"=-\dfrac{7}{2}\ln |\cos x|+\dfrac{7}{2}\ln |\sin x|+\dfrac{2}{\cos x}+C"